In an isosceles right angled triangle, the square of the hypotenuse is 98cm^2.find the length of the congruent sides?
Answer by Alan3354(69443)

In an isosceles right angled triangle, the square of the hypotenuse is 98cm^2.find the length of the congruent sides?
-------------
The sides are equal length = s
---
s^2 + s^2 = 98
2s^2 = 98
s^2 = 49
s = 7 cm